# Sparse Extended Mean-Variance-CVaR Portfolios with Short-selling

### v1

@@ 2.2.3 Subproblem of $z$
First, note that if $\lambda_{2}\delta=0,$ then we have:
$\min_{z\in\mathbb{R}^{n}}$ $\|x-z\|_{2}^{2}\qquad\textrm{subject to}\qquad L\leq z\leq U.$
In this case ${z_{*}}_{i}=U_{i}$ if $x_{i}>U_{i}$, ${z_{*}}_{i}=L_{i}$ if $x_{i}<L_{i}$, and otherwise ${z_{*}}_{i}=x_{i}$, for any $i\in[n]$, which can be rewritten as
$z_{*}=\max(L,\min(x,U)).$ (10)
Next, suppose that $\lambda_{2}\delta\neq 0$, due to the separability property of ($P_{z}$), letting $\tilde{z_{i}}=z_{i}-\phi_{i},\tilde{x}_{i}=x_{i}-\phi_{i},\tilde{L}_{i}=L_{i}-\phi_{i},$ and $\tilde{U}_{i}=U_{i}-\phi_{i}$, for $\kappa={(\lambda_{2}\delta)}/{(2\rho)}$, we shall focus on the following:
$\min_{\tilde{z}_{i}\in\mathbb{R}}\quad\frac{1}{2}(\tilde{z}_{i}-\tilde{x}_{i})^{2}+\kappa|\tilde{z}_{i}|\qquad\textrm{subject to}\qquad\tilde{L}_{i}\leq\tilde{z}_{i}\leq\tilde{U}_{i}.$
Let us assume that $\tilde{z}_{i}>0$, then we have $\tilde{z}_{i}=\tilde{x}_{i}-\kappa$. Thus, if $\tilde{x}_{i}-\kappa>0$ and $\tilde{L}_{i}\leq\tilde{x}_{i}-\kappa\leq\tilde{U}_{i}$, then $\tilde{z}_{*i}=x_{i}-\kappa$. If $0<\tilde{x}_{i}-\kappa<\tilde{L}_{i}$, we must have $\tilde{z}_{*i}=\tilde{L}_{i}$. If $\tilde{x}_{i}-\kappa>\max(\tilde{U}_{i},0)$, then $\tilde{z}_{*i}=\tilde{U}_{i}$. In the case of $\tilde{z}_{i}<0$, taking derivative gives $\tilde{z}_{i}=\tilde{x}_{i}+\kappa$. Thus, if $\tilde{x}_{i}+\kappa<0$, and $\tilde{L}_{i}\leq x_{i}+\kappa\leq\tilde{U}_{i}$, we have $z_{*i}=x_{i}+\kappa$. If $\tilde{x}_{i}+\kappa<0$, and $\tilde{x}_{i}+\kappa<L_{i}$, then we must have $\tilde{z}_{*i}=L_{i}$. Otherwise, $\tilde{x}_{i}+\kappa<0,$ and $\tilde{x}_{i}+\kappa>\tilde{U}_{i}$, so we have $\tilde{z}_{*i}=\tilde{U}_{i}$. For $\tilde{z}_{i}=0,$ we also consider three different possibilities. In this case, optimality conditions are: $-\tilde{x}_{i}+\kappa\partial(0)-T_{1}+T_{2}\in 0,\,0\leq T_{1}\perp(\tilde{L}_{i}-\tilde{z}_{i})\leq 0,\,\text{and}\,0\leq T_{2}\perp(\tilde{z}_{i}-\tilde{U}_{i})\leq 0$. It can be shown that we must have $|\tilde{x}_{i}|\leq\kappa,$ and next depending on three scenarios that $0<\tilde{L}_{i},0\in[\tilde{L}_{i},\tilde{U}_{i}],$ or $0>\tilde{U}_{i}$, we get solutions reported in below:
$\tilde{z}_{*i}:=\left\{\begin{array}[]{ll}\tilde{x}_{i}-\kappa&\qquad\mbox{if \quad$\tilde{L}_{i}\leq\tilde{x}_{i}-\kappa\leq\tilde{U}_{i}$}\qquad\text{and}\qquad\mbox{$\tilde{x}_{i}-\kappa>0$};\\ \tilde{x}_{i}+\kappa&\qquad\mbox{if \quad$\tilde{L}_{i}\leq\tilde{x}_{i}+\kappa\leq\tilde{U}_{i}$}\qquad\text{and}\qquad\mbox{$\tilde{x}_{i}+\kappa<0$};\\ L_{i}&\qquad\mbox{if \quad$0<\tilde{x}_{i}-\kappa<\tilde{L}_{i}$}\qquad\text{or}\qquad\mbox{$\tilde{x}_{i}+\kappa<\min(0,\tilde{L}_{i})$};\\ U_{i}&\qquad\mbox{if \quad$\tilde{x}_{i}-\kappa>\max(\tilde{U}_{i},0)$}\qquad\text{or}\qquad\mbox{$\tilde{U}_{i}<\tilde{x}_{i}+\kappa<0$};\\ 0&\qquad\mbox{if \quad$-\kappa\leq\tilde{x}_{i}\leq\kappa$}\qquad\text{and}\qquad\mbox{$\tilde{L}_{i}\leq 0\leq\tilde{U}_{i}$};\\ \tilde{L}_{i}&\qquad\mbox{if \quad$-\kappa\leq\tilde{x}_{i}\leq\kappa$}\qquad\text{and}\qquad\mbox{$0<\tilde{L}_{i}$};\\ \tilde{U}_{i}&\qquad\mbox{if \quad$-\kappa\leq\tilde{x}_{i}\leq\kappa$}\qquad\text{and}\qquad\mbox{$0>\tilde{U}_{i}$};\end{array}\right.$
If we summarize the above based on the original variables, we get the following:
$z_{*i}:=\left\{\begin{array}[]{ll}x_{i}-\kappa&\qquad\mbox{if \quad$L_{i}\leq x_{i}-\kappa\leq U_{i}$}\qquad\text{and}\qquad\mbox{$x_{i}-\phi_{i}-\kappa>0$};\\ x_{i}+\kappa&\qquad\mbox{if \quad$L_{i}\leq x_{i}+\kappa\leq U_{i}$}\qquad\text{and}\qquad\mbox{$x_{i}-\phi_{i}+\kappa<0$};\\ 0&\qquad\mbox{if \quad$|x_{i}-\phi_{i}|\leq\kappa$}\qquad\text{and}\qquad\mbox{$\phi_{i}\in[L_{i},U_{i}]$};\\ L_{i}&\qquad\mbox{if}\quad\left\{\begin{array}[]{ll}\phi_{i}<x_{i}-\kappa<L_{i},\qquad\mbox{or}&\\ x_{i}+\kappa<\min(\phi_{i},L_{i}),\qquad\mbox{or}&\\ |x_{i}-\phi_{i}|\leq\kappa\qquad\mbox{and}\qquad L_{i}>\phi_{i};\end{array}\right.\\ U_{i}&\qquad\mbox{if}\quad\left\{\begin{array}[]{ll}U_{i}<x_{i}+\kappa<\phi_{i},\qquad\mbox{or}&\\ x_{i}-\kappa>\max(U_{i},\phi_{i})\qquad\mbox{or}&\\ |x_{i}-\phi_{i}|\leq\kappa\qquad\mbox{and}\qquad\phi_{i}>U_{i}.\end{array}\right.\end{array}\right.$ (11)
Even though we obtained the closed-form solution of ($P_{z}$ by analysis, we will need its KKT conditions later for the convergence analysis:
$2\rho(z_{*}-x)+\lambda_{2}\delta\partial\left(\|\cdot\|_{1})\right|_{z_{*}-\phi}-T_{1}+T_{2}=0,\quad 0\leq T_{1}\perp(z_{*}-L)\geq 0,\quad\mbox{and}\quad 0\leq T_{2}\perp(U-z_{*})\geq 0$ (12)
@@ 2.2.4 Subproblem of $w$ and $\gamma$
$\min_{(w,\gamma)\in\mathbb{R}^{n}\times\mathbb{R}}$ $\ c(w,\gamma):=\lambda_{3}(\gamma+\frac{1}{m(1-\beta)}\sum_{j=1}^{m}(-d_{j}^{T}w-\gamma)^{+})+\rho\|x-w\|_{2}^{2}$ ($P_{(w,\gamma)}$)
@@
Let $\lambda_{3}\in(0,1]$ (or equivaneltly $\lambda_{1}\geq 0,\lambda_{2}\geq 0$ with $\lambda_{1}+\lambda_{2}<1$) and $\beta\in(0,1)$. Then, the problem ($P_{(w,\gamma)}$) has a unique solution.
Proof.
We first show the existence of a solution. It is enough to show that the objective function is bounded below over its feasible set. Since $\beta\in(0,1)$, we have $01-{1\over 1-\beta}<$ 0 such that
$\frac{c(w,\gamma)-\rho\|x-w\|_{2}^{2}}{\lambda_{3}}$ $=$ $\gamma+{1\over m(1-\beta)}\sum_{j=1}^{m}\max(0,-w^{T}d_{j}-\gamma)$
$=$ $\gamma+\sum_{j=1}^{m}\max(0,-{1\over m(1-\beta)}w^{T}d_{j}-{1\over m(1-\beta)}\gamma)$
$=$ $\sum_{j=1}^{m}\max({\gamma\over m},-{1\over m(1-\beta)}w^{T}d_{j}+{\gamma\over m}(1-{1\over 1-\beta}))$
$\geq$ $\sum_{j=1}^{m}{1\over{\beta\over 1-\beta}}(-{1\over m(1-\beta)}d_{j}^{T}w)$
$=$ $-{1\over m\beta}(\sum_{j=1}^{m}d_{j})^{T}w$
$\geq$ $-\frac{1}{m\beta}\|\sum_{j=1}^{m}d_{j}\|\|w\|$
$\geq$ $-\frac{1}{m\beta}\|U\|\|\sum_{j=1}^{m}d_{j}\|,$
where we used that $\max(t,u-at)\geq{u\over 1+a}$ whenever $a>0$ for $t=\gamma/m$ and $a=\beta/(1-\beta)$, and $u=-d_{j}^{T}w/m(1-\beta)$ in (2.2.4). This implies that
$c(w,\gamma)$$\geq$ $-\frac{\lambda_{3}}{m\beta}\|U\|\|\sum_{j=1}^{m}d_{j}\|+\rho\|x-w\|_{2}^{2}$ (14)
$\geq$ $-\frac{\lambda_{3}}{m\beta}\|U\|\|\sum_{j=1}^{m}d_{j}\|,$
which establishes the problem ($P_{(w,\gamma)}$) is bounded below and thus it has a solution.
Note that $w_{*}$ is unique because the Hessian of the objective function is positive definite with respect to $w$. Once $w_{*}$ is known, it can be shown that $\gamma_{*}$ is either zero or it must be equal to $-w_{*}d_{j_{*}}$ for some $j_{*}\in[m]$. To see this, without loss of generality, suppose for $j=1,\dots,\bar{m}$ with $\bar{m}\leq m$, we have $-w_{*}^{T}d_{j}-\gamma>0$, thus ($P_{(w,\gamma)}$) reduces to
$\min_{\gamma\in\mathbb{R}}$ $\gamma+\frac{1}{m(1-\beta)}\sum_{j=1}^{\bar{m}}(-w_{*}^{T}d_{j}-\gamma)=\min_{\gamma\in\mathbb{R}}\quad$ $\frac{m(1-\beta)-\bar{m}}{m(1-\beta)}\gamma.$
If $m(1-\beta)-\bar{m}\geq 0$, then $\gamma_{*}=0$, otherwise we have $\gamma_{*}=\min_{j\in[\bar{m}]}-w_{*}^{T}d_{j}$. ∎
We need the KKT conditions for $w_{*}$ in our convergence analysis later:
$\frac{-\lambda_{3}}{m(1-\beta)}\sum_{j=1}^{m}d_{j}\partial\left(.)^{+}\right|_{-d_{j}^{T}w_{*}-\gamma}+2\rho(w_{*}-x)=0,\quad\textbf{e}^{T}w_{*}=1,\quad\mbox{and}\quad L\leq w_{*}\leq U.$ (15)
Eventually, we point out that solving this convex problem can be done using a standard solver because it can be transformed into the following convex quadratic program with linear constraints:
@@ 3 Convergence Analysis
In this section, we begin by examining Algorithm 1 to address the penalty subproblem ($P_{x,y,z,w,\gamma}$) with a fixed $\rho\geq 1$. We demonstrate its efficacy in locating a saddle point for this subproblem. Subsequently, our attention shifts to Algorithm 2, where we illustrate its capability in identifying a convergent subsequence that converges a local minimizer of the original problem ($P$).
@@ 3.1 Analysis of Algorithm 1
Let $\lambda_{1}$ and $\lambda_{2}\geq 0,$ with $\lambda_{3}=1-\lambda_{1}-\lambda_{2}>0,\beta\in(0,1),\delta\geq 0$, $\eta,L,U\in\mathbb{R}^{n}$ with $L<U$ element-wise, $d_{j}\in\mathbb{R}^{n}$ for each $j\in[m],A\succeq 0,$ and $\rho\geq 1$. Consider the iterates of Algorithm 1. Then, we have
$\max\{\|x_{l}\|,\|y_{l}\|,\|z_{l}\|,\|w_{l}\|,|\gamma_{l}|\}\leq C,$ (17)
where $C>0$ is a function of the parameters of ($P$) except from $\rho$.
@@
For simplicity, we drop the subscript $l$ whenever it is clear. First, simply $\|w\|\leq\max(\|L\|,\|U\|)$. In Lemma 2.2, we also proved that $\gamma=-w^{T}d_{j}$ for some $j\in[m]$. Thus, we have $|\gamma|\leq\|w\|\max_{j\in[m]}\|d_{j}\|\leq\max(\|L\|,\|U\|)\max_{j\in[m]}\|d_{j}\|$, which implies that
$\max\{|\gamma|,\|w\|\}\leq\max(\|L\|,\|U\|)\max(1,\max_{j\in[m]}\|d_{j}\|):=\bar{C}$ (18)
Also, clearly $\|z\|\leq\max(\|L\|,\|U\|)$. Since $\|y\|\leq\|x\|$ from (2.1), we are remained to prove that $\|x\|$ is bounded. Recall the definitions in (7) and let $\alpha=0.5\textbf{e}^{T}B\textbf{e}$. We see
$\|B^{-1}\|=\frac{1}{{\lambda_{\min}(B)}}=\frac{1}{{\lambda_{1}\lambda_{\min}(A)+3\rho}}\leq\frac{1}{3{\rho}},$
and
$2\alpha=\textbf{e}^{T}B^{-1}\textbf{e}\geq\|\textbf{e}\|^{2}\lambda_{\min}(B^{-1})=\|\textbf{e}\|^{2}\lambda^{-1}_{\max}(B)=\frac{n}{\lambda_{\max}(\lambda_{1}A+3\rho)}=\frac{n}{\lambda_{1}\lambda_{\max}(A)+3\rho}.$
Also,
$\|\vartheta\|\leq 3\rho(\|y\|+\|z\|+\|w\|)+\lambda_{2}(\|\mu\|+r_{c}\|h\|)\leq 3\rho(\|x\|+2\max(\|L\|,\|U\|))+\lambda_{2}(\|\mu\|+r_{c}\|h\|)\leq 3\rho\|x\|+\mathcal{O}(1)$
Thus, the equation (8) together with $\|y\|\leq\|x\|$ leads to
$\|x\|$ $\leq\frac{1}{2}\|B^{-1}\|\left(\|\vartheta\|+\frac{1+0.5|\textbf{e}^{T}B^{-1}\vartheta|}{\alpha}\|\textbf{e}\|\right)$
$\leq\frac{1}{2}\|B^{-1}\|\left(\|\vartheta\|+\frac{1+0.5\|\textbf{e}\|\|B^{-1}\|\|\vartheta\|}{\alpha}\|\textbf{e}\|\right)$
$\leq\frac{1}{2}\|B^{-1}\|\left(\|\vartheta\|+\frac{1+0.5\|\textbf{e}\|\|\vartheta\|}{6n\rho}(\lambda_{1}\lambda_{\max}(A)+3\rho)\|\textbf{e}\|\right)$
$\leq\frac{1}{36\rho^{2}}((7.5\rho+0.5\lambda_{1}\lambda_{\max}(A))\|\vartheta\|+(\lambda_{1}\lambda_{\max}(A)+3\rho)\|\textbf{e}\|)$
$\leq\frac{1}{36\rho^{2}}(22.5\rho^{2}\|x\|+\mathcal{O}(\rho)).$
Thus, whenever $\rho\geq 1$, we have $\|x\|\leq{\mathcal{O}(1)}/{13.5\rho}\leq\mathcal{O}(1).$ Consequently, we can say that $\max\{\|x_{l}\|,$ $\|y_{l}\|,\|z_{l}\|,\|w_{l}\|,|\gamma_{l}|\}$ is bounded above and its bound is independent from $\rho$. ∎
The provided lemma above proves that any sequence formed by Algorithm 1 is bounded and specifically, this bound is independent of $\rho$ whenever $\rho\geq 1$, which is the case for our penalty decomposition algorithm. Hence, every sequence produced by Algorithm 1 possesses at least one accumulation point. The subsequent theorem further confirms that each accumulation point is indeed a saddle point of ($P_{x,y,z,w,\gamma}$).
Theorem 3.1.
Let $\{(x_{l},y_{l},z_{l},w_{l},\gamma_{l})\}$ be a sequence generated by Algorithm 1 for solving ($P_{x,y,z,w,\gamma}$). Then, its accumulation point $(x_{*},y_{*},z_{*},w_{*},\gamma_{*})$ is a saddle point of the nonconvex problem ($P_{x,y,z,w,\gamma}$). Moreover, $\left\{q_{\rho}(x_{l},y_{l},z_{l},w_{l},\gamma_{l})\right\}$ is a non-increasing sequence.
Proof.
By observing definitions of $x_{l+1},y_{l+1},z_{l+1},$ and $(w_{l+1},\gamma_{l+1})$ in steps 6-9 of Algorithm 1, we get
$q_{\rho}(x_{l+1},y_{l+1},z_{l+1},w_{l+1},\gamma_{l+1})$$\leq$ $q_{\rho}(x_{l+1},y_{l+1},z_{l+1},w,\gamma);\qquad\forall w\in\mathcal{W}\quad\mbox{and}\quad\forall\gamma\in\mathbb{R},$
@@
This leads to the following:
$q_{\rho}(x_{l+1},y_{l+1},z_{l+1},w_{l+1},\gamma_{l+1})$ $\leq q_{\rho}(x_{l+1},y_{l+1},z_{l+1},w_{l},\gamma_{l})$ (20)
@@
$\leq q_{\rho}(x_{l},y_{l},z_{l},w_{l},\gamma_{l});\quad\forall l\in\mathbb{N},$
which shows that $q_{\rho}(x_{l},y_{l},z_{l},w_{l},\gamma_{l})$ is non-increasing sequence. Since it is also bounded below when $(x_{l},y_{l},z_{l},w_{l},\gamma_{l})\in\mathcal{X}\times\mathcal{Y}\times\mathcal{Z}\times\mathcal{W}\times\mathbb{R}$, we conclude its convergence.
Next, let $(x_{*},y_{*},z_{*},w_{*},\gamma_{*})$ be an accumulation point of $\{(x_{l},y_{l},z_{l},w_{l},\gamma_{l})\}$, there exists a subsequence $\bar{\mathcal{L}}$ such that $\lim_{l\in\bar{\mathcal{L}}\to\infty}(x_{l},y_{l},z_{l},w_{l},\gamma_{l})=(x_{*},y_{*},z_{*},w_{*},\gamma_{*})$. The continuity of $q_{\rho}(x_{l},y_{l},z_{l},w_{l},\gamma_{l})$ yields
@@
By the continuity of $q_{\rho}(x_{l},y_{l},z_{l})$ and taking the limit of both sides of (3.1) as $l\in\bar{\mathcal{L}}\to\infty$ , we have
$q_{\rho}(x_{*},y_{*},z_{*},w_{*},\gamma_{*})$$\leq$ $q_{\rho}(x_{*},y_{*},z_{*},w,\gamma);\qquad\forall x\in\mathcal{W}\quad\mbox{and}\quad\forall\gamma\in\mathbb{R},$
$q_{\rho}(x_{*},y_{*},z_{*},w_{*},\gamma_{*})$$\leq$ $q_{\rho}(x,y_{*},z,w_{*},\gamma_{*});\qquad\forall z\in\mathcal{Z},$

### v2

@@ 2.2.3 Subproblem of $z$
Lemma 2.3.
For $i\in[n]$, the solution to the problem ($P_{z}$) is as follows:
[…]
where $\kappa=\lambda_{2}\delta/\rho$.
Even though we obtained the closed-form solution of ($P_{z}$) by analysis, we mention that its KKT conditions are:
$2\rho(z_{*}-x)+\lambda_{2}\delta\partial\left(\|\cdot\|_{1})\right|_{z_{*}-\phi}-T_{1}+T_{2}=0,\quad 0\leq T_{1}\perp(z_{*}-L)\geq 0,\quad\mbox{and}\quad 0\leq T_{2}\perp(U-z_{*})\geq 0.$
@@ 2.2.4 Subproblem of $w$ and $\gamma$
$\min_{(w,\gamma)\in\mathbb{R}^{n}\times\mathbb{R}}$ $\ c(w,\gamma)=\lambda_{3}(\gamma+\frac{1}{m(1-\beta)}\sum_{j=1}^{m}(-d_{j}^{T}w-\gamma)^{+})+\rho\|x-w\|_{2}^{2}$ ($P_{(w,\gamma)}$)
@@
Let $\lambda_{3}\in(0,1]$ (or equivalently $\lambda_{1}\geq 0,\lambda_{2}\geq 0$ with $\lambda_{1}+\lambda_{2}<1$) and $\beta\in(0,1)$. Then, the problem ($P_{(w,\gamma)}$) has a unique solution.
The KKT conditions for $w_{*}$ are:
$\frac{-\lambda_{3}}{m(1-\beta)}\sum_{j=1}^{m}d_{j}\partial\left(.)^{+}\right|_{-d_{j}^{T}w_{*}-\gamma}+2\rho(w_{*}-x)=0,\quad\textbf{e}^{T}w_{*}=1,\quad\mbox{and}\quad L\leq w_{*}\leq U.$
Eventually, we point out that this convex problem can be solved using a standard solver because it can be transformed into the following convex quadratic program with linear constraints:
@@ 3 Convergence Analysis
In this section, we begin by examining Algorithm 1 to address the penalty subproblem ($P_{x,y,z,w,\gamma}$) with a fixed $\rho\geq 1$. We demonstrate its efficacy in locating a saddle point for this subproblem. Subsequently, our attention shifts to Algorithm 2, where we illustrate its capability in identifying a convergent subsequence that converges to a Lu–Zhang minimizer of the original problem ($P$).
@@ 3.1 Analysis of Algorithm 1
Let $\lambda_{1}$ and $\lambda_{2}\geq 0,$ with $\lambda_{3}=1-\lambda_{1}-\lambda_{2}\geq 0,\beta\in(0,1),\delta\geq 0$, $\eta,L,U\in\mathbb{R}^{n}$ with $L<U$ element-wise, $d_{j}\in\mathbb{R}^{n}$ for each $j\in[m],A\succeq 0,$ and $\rho\geq 1$. Consider the iterates of Algorithm 1. Then, we have
$\max\{\|x_{l}\|,\|y_{l}\|,\|z_{l}\|,\|w_{l}\|,|\gamma_{l}|\}\leq C,$
where $C>0$ is a function of the parameters of ($P$), except for $\rho$.
This lemma proves that any sequence formed by Algorithm 1 is bounded; specifically, this bound is independent of $\rho$ whenever $\rho\geq 1$, which is the case for our penalty decomposition algorithm. Hence, every sequence produced by Algorithm 1 possesses at least one accumulation point. The next theorem further confirms that each accumulation point is a saddle point of ($P_{x,y,z,w,\gamma}$).
We note that a saddle point (block-coordinate minimizer) refers to a point $(x_{*},y_{*},z_{*},w_{*},\gamma_{*})$ that minimizes $q_{\rho}$ with respect to each block separately while keeping the other blocks fixed (see (10)). This notion is generally weaker than full first-order stationarity or the KKT conditions for the joint problem. If $q_{\rho}$ is continuously differentiable and each block-feasible set $\mathcal{X},\mathcal{Y},\mathcal{Z},\mathcal{W}$ is closed and convex, then (10) is equivalent to the first-order optimality condition
$0\in\nabla q_{\rho}(x_{*},y_{*},z_{*},w_{*},\gamma_{*})\;+\;N_{\mathcal{X}\times\mathcal{Y}\times\mathcal{Z}\times\mathcal{W}\times\mathbb{R}}(x_{*},y_{*},z_{*},w_{*},\gamma_{*}),$
where $N_{C}(u)$ denotes the (convex) normal cone of a closed convex set $C$ at $u$, so the limit point is (Clarke/KKT) stationary. In our setting, however, $\mathcal{Y}$ is nonconvex due to the cardinality constraint, so we do not claim full stationarity; Theorem 3.1 asserts convergence to a saddle point (block-coordinate minimizer).
Theorem 3.1.
Let $\{(x_{l},y_{l},z_{l},w_{l},\gamma_{l})\}$ be a sequence generated by Algorithm 1 for solving ($P_{x,y,z,w,\gamma}$). First, $\left\{q_{\rho}(x_{l},y_{l},z_{l},w_{l},\gamma_{l})\right\}$ is a non-increasing sequence. Second, any accumulation point $(x_{*},y_{*},z_{*},w_{*},\gamma_{*})$ of $\{(x_{l},y_{l},z_{l},w_{l},\gamma_{l})\}$ is a saddle point of the nonconvex problem ($P_{x,y,z,w,\gamma}$), that is,
$x_{*}$ $\in\arg\min_{x\in\mathcal{X}}q_{\rho}(x,y_{*},z_{*},w_{*},\gamma_{*}),$ (10)
$y_{*}$ $\in\arg\min_{y\in\mathcal{Y}}q_{\rho}(x_{*},y,z_{*},w_{*},\gamma_{*}),$
$z_{*}$ $\in\arg\min_{z\in\mathcal{Z}}q_{\rho}(x_{*},y_{*},z,w_{*},\gamma_{*}),$
$(w_{*},\gamma_{*})$ $\in\arg\min_{\begin{subarray}{c}w\in\mathcal{W}\\ \gamma\in\mathbb{R}\end{subarray}}q_{\rho}(x_{*},y_{*},z_{*},w,\gamma).$
@@
By observing the definitions of $x_{l+1},y_{l+1},z_{l+1},$ and $(w_{l+1},\gamma_{l+1})$ in steps 6–9 of Algorithm 1, we obtain
$q_{\rho}(x_{l+1},y_{l+1},z_{l+1},w_{l+1},\gamma_{l+1})$ $\leq$ $q_{\rho}(x_{l+1},y_{l+1},z_{l+1},w,\gamma);\qquad\forall w\in\mathcal{W}\ \mbox{and}\ \forall\gamma\in\mathbb{R},$
@@
This leads to the following ($\forall l\in\mathbb{N}$):
$q_{\rho}(x_{l+1},y_{l+1},z_{l+1},w_{l+1},\gamma_{l+1})$ $\leq q_{\rho}(x_{l+1},y_{l+1},z_{l+1},w_{l},\gamma_{l})$
@@
$\leq q_{\rho}(x_{l},y_{l},z_{l},w_{l},\gamma_{l}),$
which shows that $q_{\rho}(x_{l},y_{l},z_{l},w_{l},\gamma_{l})$ is a non-increasing sequence. Furthermore, by Lemma 3.1, the iterates $\{(x_{l},y_{l},z_{l},w_{l},\gamma_{l})\}_{l\in\mathbb{N}}$ remain in a compact set. Since $q_{\rho}$ is continuous, the sequence of objective values $\{q_{\rho}(x_{l},y_{l},z_{l},w_{l},\gamma_{l})\}_{l\in\mathbb{N}}$ is bounded below. Together with the fact established earlier that $\{q_{\rho}(x_{l},y_{l},z_{l},w_{l},\gamma_{l})\}_{l\in\mathbb{N}}$ is non-increasing, it follows that this sequence converges.
Next, let $(x_{*},y_{*},z_{*},w_{*},\gamma_{*})$ be an accumulation point of $\{(x_{l},y_{l},z_{l},w_{l},\gamma_{l})\}_{l\in\mathbb{N}}$; i.e., along some infinite index set $\bar{\mathcal{L}}\subset\mathbb{N}$ we have $(x_{l},y_{l},z_{l},w_{l},\gamma_{l})\to(x_{*},y_{*},z_{*},w_{*},\gamma_{*})$. Since $\mathcal{X},\mathcal{Y},\mathcal{Z},$ and $\mathcal{W}$ are closed, the limit is feasible. Because $\bar{\mathcal{L}}$ is infinite, the shifted set $\bar{\mathcal{L}}-1:=\{\,l-1:\ l\in\bar{\mathcal{L}},\,l\geq 1\,\}$ is also infinite. Replacing $l$ by l1l-1 in (3.1) and letting $l\to\infty$ along $\bar{\mathcal{L}}$, continuity of $q_{\rho}$ yields:
@@
Taking limits in (3.1) along $\bar{\mathcal{L}}$ and using continuity of $q_{\rho}$ gives
$q_{\rho}(x_{*},y_{*},z_{*},w_{*},\gamma_{*})$ $\leq$ $q_{\rho}(x_{*},y_{*},z_{*},w,\gamma);\qquad\forall w\in\mathcal{W}\ \mbox{and}\ \forall\gamma\in\mathbb{R},$
$q_{\rho}(x_{*},y_{*},z_{*},w_{*},\gamma_{*})$ $\leq$ $q_{\rho}(x_{*},y_{*},z,w_{*},\gamma_{*});\qquad\forall z\in\mathcal{Z},$
@@
Equivalently, (10) holds. Thus, $(x_{*},y_{*},z_{*},w_{*},\gamma_{*})$ is a saddle point (block-coordinate minimizer) of ($P_{x,y,z,w,\gamma}$). ∎